It is claimed that the average weight of a bag of biscuits is 250 grams with a standard deviation of 20.5 grams Would you agree to this claim if a random


1
Expert's answer
2021-06-01T17:47:59-0400

Would you agree to this claim if random sample of 50 bags of biscuits showed an average weight of 240 grams, using a 0.05 level of significance? 


Null Hypothesis "H_0:\mu=250"

Alternate Hypothesis "H_1:\mu\neq 250"


Standard deviation "\sigma =20.5"

The population standard deviation is known and the sample size "\ud835\udc5b" is large "(\ud835\udc5b \u2265 30)"


Hence, the test static is "z=\dfrac{\bar x-\mu}{\sigma/\sqrt n}" and


the value is "z=\dfrac{\bar x-\mu}{\sigma/\sqrt n}=\dfrac{240-255}{20.5/\sqrt{50}}=-5.174"


The test is two tailed (because "\ud835\udc3b_1" contains inequality "\u2260" ) and the critical values are "\ud835\udc67_{\u22120.025} = \u22121.96" and "\ud835\udc67_{0.025} = 1.96" . Since "\ud835\udc67 < \u2212\ud835\udc67_{0.025}" , i.e. the "\ud835\udc67" lies to the left of the critical value, we reject the null hypothesis. 





We would not agree with a claim, the true average weight of a bag of biscuits is not 250 grams. Answer: the true average weight of a bag of biscuits is not 250 grams


Answer: The true average weight of a bag of biscuit is not 250 grams.